A system has a net charge of −5.328×10−17 C . It consists of electrons and protons, in a total number of 1505 particles.Problem: 19.12 A system has a net charge of -5.328x10 17 C. It consists of electrons and protons, in a total number of 1505 particles. Part A What is the number of electrons in the system? Submit My Answers Give Up incorrect; Try Again Part B What is the mass of this system? Submit My Answers Give Up kg
Answer #1

Let the no.of protons = a

no.of electrons=b

Total no.of particles, a+b= 1505  .............(1)

Net charge, Q=−5.328×10^−17 C

But net charge id due to protons and electrons

Q= ae-be

where e is charge on electron or proton

Therefore, Q=(a-b)e=(a-b)*1.6*10^-19

but  Q=−5.328×10^−17 C =(a-b)*1.6*10^-19

a-b=3.33*10^2

a-b=-333 .............(2)

solving (1) and (2) we get

The no.of protons, a= 586

b=919The no.of electrons, b=919

Answer (b)

Mass of the system, M=a*m1+b*m2

where m1 is mass of proton = 1.9*10^-27 kg

m2 is mass of electron=9.3^-31 kg

M=586*1.9*10^-27 kg + 919*9.3^-31 kg

M=1113.4*10^-27+0.85*10^-27

M=1114.25*10^-27 kg